OpenStudy (anonymous):

Maria gave 1/2 of her jellyjeans to Carole. Carole gave 1/3 of those to Austin. Austin gave 1/4 of those to Tony. If Tony recieved 2 jelly beans, how many did Maria start with?

OpenStudy (gorv):

LET MARIA HAD X JEELY

OpenStudy (gorv):

carole=x/2 austin=1/3 of carole=1/3*()x/2)==x/6 tony= 1/4 of x/6=x/24 tony have=2=x/24 solve it now